OpenStudy (anonymous):

Part 1: Show the work to solve |3x – 2| < 7 Part 2: Describe the graph of the solution in words.

OpenStudy (cwrw238):

solve the 2 equations 3x - 2 < 7, 3x - 2 > -7

OpenStudy (anonymous):

What

OpenStudy (anonymous):

K

OpenStudy (austinl):

Solve \(3x-2<7\) for x.

OpenStudy (cwrw238):

the above equations are equivalent to |3x - 2| < 7

OpenStudy (anonymous):

So do you add 7 to each side then divide by what

OpenStudy (austinl):

No, you would add two to both sides.

OpenStudy (cwrw238):

first do what austini says and combine with solution of other equation

OpenStudy (anonymous):

Cool

OpenStudy (cwrw238):

no you need to get rid of the -2 so you add 2 to each side

OpenStudy (anonymous):

Alright to the 7 an 2 add 2

OpenStudy (anonymous):

So you marked the two 2s out an add 7+2 which is 9

OpenStudy (anonymous):

Them what you divide by

OpenStudy (cwrw238):

3x - 2 + 2 = 7 + 2 3x = 9 divide by 3

OpenStudy (anonymous):

So divide both by 3

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

So z equals 9 ten divide them by 3

OpenStudy (anonymous):

An 2z =a 9?

OpenStudy (cwrw238):

that gives you the value of x

OpenStudy (anonymous):

I put theslash down to separate the two things what do I need to put below the line 2z it's 9

OpenStudy (anonymous):

An say if you have all C's an one D what grade point adverage what would that be that's off topic ok but was just curious